SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the Trigger Operation Meeting (July 5, 2005)

Attendees: Prafulla Behera, Kris Chaisanguanthum, Al Eisner, Fang Fang, Larry Gladney, Selina Li, Brian Lindquist, Dong Su, Dan Walker.

Operations

  • The monthly Trigger status report was given last Wednesday. Three questions were asked.
    1. Any progress on the GLT deadtime reduction?
    2. Do we have any plan on working on solving the problem where the DCZ crates have to be power-cycled when the master crate lost the timing info? For example, when the ROM was power-cycled, the TSF crate has to be power cycled afterward, an issue with the digital clock manager.
    3. How much longer do we plan to run the parasitic DCT? Selina replied "till August".
  • DIRC timing problem (see discussion under "GLT" below)
  • TSF out-of-frame error incidence happened again on Wednesday June 29. (see "DCZ" below)
  • ZPD damaged events happened yesterday on July 4. (see "DCZ" below)

EMT

  • The EMT calibration had problem with serial number being swapped in some TPC cables, which needs updating.
  • Dan showed some plots for mu pair events as well as bhabha phi energy. It was suggested that the entries that showed up below threshold could be background such as gamma from gamma mu+mu-.

DCT/DCZ

  • The TSF-X15 had out-of-frame error again last Wednesday. See more detailed in HN posted. Su Dong suggested to swap the X15 with another slot to see if the problem will follow the board. This swapping will be done tomorrow during MD.

GLT

  • DIRC saw a timing issue involving trigger lines (see more discussion in the associated HN posted). Al also saw some shift in SVT for the first 10K events or so when he ran a playback job on run 55662, but this effect got smeared out with more statistics. We wonder same thing happened to DIRC too if they look at the result from a playback job.

Level 3

  • No L3 Node trim for 2 weeks now.
  • Prafulla updated the trigConfig for June.

Data Quality and Monitoring

  • The DQM shifter posted a HN to inform us some entries were seen in the EMT-GLT and DCT-GLT transmission error plots. But we don't see it. Need more clarification.

This page is created by Selina Li (selina@slac.stanford.edu); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.